# Modeling and Analysis of Land-to-Ship Maritime Wireless Channels at 5.8 GHz

### v1

@@ Abstract
Maritime channel modeling is crucial for designing robust communication systems in marine environments, where factors like waves and wind impact signal propagation. This article investigates land-to-ship maritime wireless channel characteristics at 5.8 GHz based upon an extensive measurement campaign, with concurrent hydrological and meteorological information collection. First, a novel large-scale path loss model with physical foundation and high accuracy is proposed for dynamic marine environments. Then, we introduce the concept of sea-wave-induced fixed-point (SWIFT) fading, a peculiar phenomenon in maritime scenarios that captures the impact of sea surface fluctuations on received power. An enhanced two-ray model incorporating vessel rotational motion is propounded to simulate the SWIFT fading, showing good alignment with measured data, particularly for modest antenna movements. Next, the small-scale fading is studied by leveraging a variety of models including the two-wave with diffuse power (TWDP) and asymmetric Laplace distributions, with the latter performing well in most cases, while TWDP better captures bimodal fading in rough seas. Furthermore, maritime channel sparsity is examined via the Gini index and Rician $K$ factor, and temporal dispersion is characterized. The resulting channel models and parameter characteristics offer valuable insights for maritime wireless system design and deployment.
Index Terms:
Maritime channel measurements, channel model, path loss, fixed-point fading, small-scale fading, channel sparsity.
@@ I Introduction
Channel modeling typically begins with measurements in various representative environments to extract key channel parameters and develop corresponding models, including standardized or widely adopted models such as WINNERII[5], 3rd Generation Partnership Project (3GPP)[6], the Simulation of Indoor Radio Channel Impulse Response (CIR) Models (SIRCIM)[7], the Simulation of Mobile Radio CIR Models (SMRCIM)[8], and NYUSIM[9]. Numerous channel measurements and models exist for 5G/6G terrestrial scenarios (e.g., [14, 11, 12, 13, 10] and references therein). However, when considering the extension of terrestrial broadband mobile communication systems to nearshore areas, the unique attributes of the maritime environment necessitate new channel measurement and modeling efforts. Unlike terrestrial scenarios, the maritime radio propagation environment is significantly affected by various factors such as wind speed, temperature, sea states, and vessel motion. Additionally, scatterers and users are usually sparsely distributed over the vast sea surface, rendering traditional assumptions like Rayleigh fading invalid. Instead, maritime wireless channels exhibit sparsity and location dependency[3]. Furthermore, sea surface fluctuations are time-varying and not confined to fixed temporal scales, leading to more complex random undulations in nearshore wireless channels [15]. Thus, accurate modeling must account for the stochastic effects of wave dynamics.
@@
Among the existing maritime channel measurements, researchers mainly focused on the S-band and C-band. Lee et al. conducted land-to-ship channel measurements in the Yellow Sea of China, analyzing the received signal strength and the small-scale fading in the line-of-sight (LoS) propagation region at 2.4 GHz, and the results indicated that small-scale fading is best characterized by the Rician distribution [16]. Authors in [17] investigated the applicability of several terrestrial wireless channel models in maritime environments at 2 GHz and found that the ITU-R model[18] showed the best agreement with the measured results. By adding a Gaussian random variable, the work [19] described the randomness introduced by environmental changes in maritime channels. The researchers in [20] conducted land-to-ship channel measurements at 3.2 GHz and analyzed how islands affect wireless propagation, revealing that cluster delays shifted linearly with distance and small-scale fading changed from Rician to log-normal distribution when the ship passed through islands. In [21], measurements were conducted at 5 GHz in tropical sea regions, and it was proposed that a third refracted path, in addition to the direct and reflected paths, may occur during long-distance maritime transmission due to the evaporation duct effect over the sea. The measurement results showed excellent agreement with the theoretical curves, indicating that considering the duct effect is necessary in long-distance maritime scenarios. In[22], channel measurements were carried out at 5.8 GHz and instantaneous power delay profiles (PDPs) of wireless links were acquired between buoys and ships under various multipath conditions, followed by the investigation on delay dispersion characteristics. In [23], 5.8 GHz fixed WiMAX measurements were conducted in harbor environments to study the impacts of multipath effects, Doppler shift, and vessel motion on wireless channels. For the same frequency band, a land-to-ship channel measurement campaign was launched [24, 25], accompanied with some preliminary analysis on path loss, small-scale fading, and root-mean-square (RMS) delay spread. Ship-to-ship measurements at 5.9 GHz with a maximum distance of 2.6 km were performed[26], where the statistical properties of several channel parameters were analyzed. The authors of [27] conducted measurements with different receiver (Rx) heights at 8 GHz in the South China Sea. Among the fitting results of various path loss models, the log-distance model exhibited the smallest root mean square error (RMSE), while the parabolic equation model provided additional information on interference nulls. For small-scale fading, the Rician distribution showed the best Kolmogorov–Smirnov (K-S) test metrics compared to other commonly used distributions.
@@
Attention is also gradually shifting towards even higher frequency bands. A beyond-the-horizon maritime channel measurement campaign was conducted at 25 GHz for fixed-to-fixed point links[28]. Based on the measurement data, the authors analyzed the short-term variations in received signal power caused by environmental changes and presented empirical models. To better align with real-world conditions, some studies have applied modifications to the traditional two-ray model to more comprehensively account for fading produced by maritime environmental factors. For instance, the work [29] combined reflection from the sea surface and antenna height variation to propose a modified two-ray model through theoretical analysis, while [30] introduced an exponential coefficient into the two-ray model formula to correct for fluctuations affected by environmental factors via ray tracing, and achieved better prediction results in the millimeter-wave frequency band.
@@
However, the existing works on maritime channel measurements and modeling either have not considered the impact of various environmental factors on channel characteristics, or the investigated distances are limited, or the channel parameters and models are not comprehensive enough. In this paper, based on real-world channel measurement data at 5.8 GHz for the land-to-ship scenario in the Yellow Sea of China [24, 25], we conduct systematic modeling and analysis on a variety of key channel parameters such as large-scale path loss, power fading on different time scales, channel sparsity, and temporal statistics. The main novelty and contributions of this work, compared with the current literature including our previous study [24] and [25], are as follows:
@@
  • A series of large-scale fading models are employed to characterize the location-dependent path loss. More importantly, a new dual-slope large-scale path loss model is propounded, which embodies a solid physical basis, high modeling accuracy, and super adaptability over a wide range of distances and sea conditions. In addition, the two-wave with diffuse power (TWDP) model and asymmetric Laplace function are utilized to depict the small-scale fading, in addition to the conventional Rician, Nakagami-m, and lognormal distributions.
@@
  • Channel sparsity for the marine propagation environment is investigated in depth, where different metrics including the Gini index and Rician $K$ factor are adopted to quantify the sparsity level. The Gini index stems from economics and has been shown to predict channel sparsity accurately, and it is garnering attention recently in research on land channels but has not been considered for maritime scenarios so far to our best knowledge.
The remainder of this paper is organized as follows. In Section II, the maritime wireless channel measurement system and procedure are presented, along with a brief description of the data processing method. In Sections III to V focus on the large-scale, SWIFT, and small-scale fading, respectively. Channel sparsity is analyzed in Section VI, followed by the temporal statistics in Section VII. Conclusions are drawn in Section VIII.
@@ II-A Measurement System
At the Rx, also as shown in Fig. 1, the signal is received by an omnidirectional antenna with a gain of 13 dBi. The received signal is then amplified by a low noise amplifier with 27.4 dB gain and subsequently processed by a spectrum and signal analyzer (R&S® FSW43) for analog-to-digital conversion and IQ demodulation. A 10 MHz GPS signal is employed at both the Tx and Rx to synchronize their respective transmission and reception times. Environmental information is collected along with the measured data. Fig. 2 presents photographs of the measurement equipment and the overall measurement scenario over the open sea. More details of the measurement system parameter settings are listed in Table I.
@@ II-B Measurement Procedure
The Tx antenna was fixed on the rooftop of a 25 m-high residential building, while the Rx antenna was mounted on the mast of a vessel’s deck as shown in Fig. 2(c). A Kestrel 5500 weather meter was installed on the vessel to record meteorological parameters such as the wind speed, air temperature, humidity, and atmospheric pressure, as displayed in Fig. 2(d). Additional marine environmental information, including sea water temperature and wave height, was recorded using the monitoring buoys in the sea area of Qidong City, as illustrated in Fig. 2(e). The measurements were carried out over two consecutive days. On the first day, the average relative wind speed over the sea surface was about 7.7 m/s, higher than the average relative wind speed of 5.6 m/s on the second day, while the other measurement conditions were almost the same. Correspondingly, there were significant differences in the sea wave conditions between the two days.
On both days, the vessel carrying the Rx sailed away from the shore along the due east direction, and fixed-point measurements were performed at more than ten predetermined Rx locations as delineated in Fig. 3, with an Rx antenna height of 4 m. On the return trip of the vessel, measurements were taken without fixed Rx locations, maintaining a constant vessel speed of 9.6 knots while continuously recording the received signal strength. The vessel’s direction was always aligned with the direction of the Tx antenna boresight, and the maximum Tx-Rx separation distance reached 33 km. On the second day, due to the decreased wind speed, additional fixed-point measurements with a lower Rx antenna height of 1 m were conducted for comparison. In the continuous-distance measurements, 14 sets of IQ data were collected within about 60 milliseconds at each measurement location, with adjacent measurement locations spaced about 0.02 km apart. While in the fixed-point measurements, 31 groups of the aforementioned 14 sets of IQ data were recorded with an approximately three-second time interval between adjacent groups, producing a total of 434 sets of IQ data at each location. More details of the measurement procedure are provided in Table II. As the measurements were conducted in winter, the sea surface temperature was around 3°C, resulting in very weak evaporation effect [3]. Therefore, the evaporation duct effect is neglected in the following analysis.
@@ II-C Data Processing
where $\otimes$ and * denote the correlation and convolution operations, respectively, $h(t,\tau)$ is the ground-truth CIR. The PDP is the squared-magnitude of the CIR, i.e., $|\hat{h}(t,\tau)|^{2}$. Examples of measured PDPs in the continuous-distance measurements are portrayed in Fig. 4.  [^3] It is noteworthy that the received power drops substantially around 4 km compared with its adjacent distances in Fig. 4, corresponding to the most dominant peak in path loss in Fig. 5(b), which will be elaborated in the next section.
@@
Based on the measured data above, a comprehensive characterization of maritime channels requires an understanding of different fading mechanisms occurring across distinct temporal scales. The instantaneous received power can be conceptually expressed as
@@ III-A Empirical Distance-Dependent Path Loss Models
$d_{\text{LoS}}^{\text{vision}}=\sqrt{h_{\text{t}}^{2}+2h_{\text{t}}r_{\text{e}}}+\sqrt{h_{\text{r}}^{2}+2h_{\text{r}}r_{\text{e}}},$ (4)
@@
$\text{PL}^{\text{two-ray}}[\text{dB}]=20\log_{10}\Bigg{\{}\frac{4\pi d}{\lambda}\bigg{/}\left[2\sin\left(\frac{2\pi h_{\text{t}}h_{\text{r}}}{\lambda d}\right)\right]\Bigg{\}},$ (5)
@@
In realistic maritime channels, the sea-surface reflection coefficient $\Gamma$ can be influenced by multiple factors, hence we develop a modified two-ray (MTR) model as follows
$\text{PL}^{\text{MTR}}[\text{dB}]=20\log_{10}\Bigg{\{}\frac{4\pi d}{\lambda}\bigg{/}\left|1+DSR\Gamma e^{-j\frac{2\pi\Delta d}{\lambda d}}\right|\Bigg{\}},$ (6)
@@
$D=\left(1+\frac{2d_{1}d_{2}}{r_{e}(h_{\text{t}}+h_{\text{r}})}\right),$ (7)
@@
$\begin{split}\Lambda(\tan\theta_{i})=&~{}\frac{1}{2}\Bigg{\{}\sqrt{\frac{2}{\pi}}\frac{\beta_{0}}{\tan\theta_{i}}\exp{\left[-\left(\frac{\tan\theta_{i}}{\sqrt{2}\beta_{0}}\right)^{2}\right]}\\ &~{}-\text{erfc}\left(\frac{\tan\theta_{i}}{\sqrt{2}\beta_{0}}\right)\Bigg{\}}.\end{split}$ (9)
@@
Figure 5: Measured path loss data and various path loss models for the two measurement days.
@@
In (12) and (13), $\text{PL}^{\text{CI}}$ and $\text{PL}_{\text{Dual}}^{\text{CI}}$ denote the path loss in dB over frequency and distance, $n$, $n_{1}$, and $n_{2}$ represent the path loss exponents (PLEs) within different distance ranges, and $X_{\sigma}^{\text{CI}}$ is a zero-mean log-normally distributed random variable with standard deviation $\sigma$ describing large-scale shadow fading (SF).
@@ III-B Measured Results and Proposed Path Loss Model
${\color[rgb]{0,0,0}\text{PL}^{\text{Dual-CI-MTR}}(f,d)[\text{dB}]\\ =\begin{cases}10n_{1}\log_{10}\left\{\frac{4\pi fd}{c\left|1+DSR\Gamma e^{-j\frac{2\pi f\Delta d}{cd}}\right|}\right\},\quad\quad\quad\quad\quad\quad\quad\quad~{}\text{for }d\leq d_{\text{break}}\\ 10n_{1}\log_{10}\left\{\frac{4\pi fd_{\text{break}}}{c\left|1+DSR\Gamma e^{-j\frac{2\pi f\Delta d}{cd_{\text{break}}}}\right|}\right\}+10n_{2}\log_{10}\left(\frac{d}{d_{\text{break}}}\right),\text{ }\text{for }d>d_{\text{break}}\end{cases}+X_{\sigma}^{\text{Dual-CI-MTR}}}$ (14)
@@
In view of the trend of the measured data and observations above, we propound a new path loss model named dual-slope CI-MTR model integrating the CI and MTR models, which is expressed in (14). In this model, the MTR model incorporating an adaptive PLE $n_{1}$ is proposed to delineate the path loss within $d_{\text{break}}$, in conjunction with a second slope $n_{2}$ similar to that in the dual-slope CI model for distances beyond $d_{\text{break}}$. The dual-slope CI-MTR model captures the oscillating behavior of the path loss within $d_{\text{break}}$, as demonstrated by the multiple peaks of the dashed blue curve in Fig. 5. Furthermore, it allows for more flexible and accurate characterization of the path loss based on environmental conditions via $n_{1}$ and $n_{2}$. The slope values in the dual-slope CI-MTR model are similar to those in the dual-slope CI model in relevant literature (e.g., around 2 and 6 for calm sea in[23]), with lower $\sigma$ (e.g., over 8 dB in the literature) probably due to improved fitting within $d_{\text{break}}$. Since no standardized maritime path loss model with PLEs exists to our best knowledge, we compare the PLEs and SF standard deviations in our dual-slope CI-MTR model against 3GPP TR 38.901’s rural macro (RMa) LoS scenario. Most PLEs align (2 and 4 for the two slopes), except for $n_{2}$ on Day 2, likely due to Earth-curvature diffraction effects absent in the 3GPP RMa’s 10 km range limit. Besides, the $\sigma$ on Day 1 is higher than the RMa scenario (4-6 dB), implying windy environments render more widely spread path loss values.

### v2

@@ Abstract
Maritime channel modeling is crucial for designing robust nearshore communication systems, yet reliable models that account for the dynamic marine environment with varying sea waves, wind conditions, and vessel motions remain scarce. This article investigates land-to-ship maritime wireless channel characteristics at 5.8 GHz based upon an extensive measurement campaign, with concurrent hydrological and meteorological information collection. First, a novel large-scale path loss model with physical foundation and high accuracy is proposed for dynamic marine environments. Then, we introduce the concept of sea-wave-induced fixed-point (SWIFT) fading, a peculiar phenomenon in maritime scenarios that captures the impact of sea surface fluctuations on received power. An enhanced two-ray model incorporating vessel rotational motion is propounded to simulate the SWIFT fading, showing good alignment with measured data, particularly for modest antenna movements. Next, the small-scale fading is studied by leveraging a variety of models including the two-wave with diffuse power (TWDP) and asymmetric Laplace distributions, with the latter performing well in most cases, while TWDP better captures bimodal fading in rough seas. Furthermore, maritime channel sparsity is examined via the Gini index and Rician $K$ factor, and temporal dispersion is characterized. The resulting channel models and parameter characteristics offer valuable insights for maritime wireless system design and deployment.
@@ I Introduction
Channel modeling typically begins with measurements in various representative environments to extract key channel parameters and develop corresponding models, including standardized or widely adopted models such as WINNER II[5], 3rd Generation Partnership Project (3GPP)[6], the Simulation of Indoor Radio Channel Impulse Response (CIR) Models (SIRCIM)[7], the Simulation of Mobile Radio CIR Models (SMRCIM)[8], and NYUSIM[9]. Numerous channel measurements and models exist for 5G/6G terrestrial scenarios, e.g., [10],[11],[12],[13],[14] and references therein, which offer insightful perspectives on terrestrial channel modeling and characteristics analysis. However, when considering the extension of terrestrial broadband mobile communication systems to nearshore areas, the unique attributes of the maritime environment necessitate new channel measurement and modeling efforts. Unlike terrestrial scenarios, the maritime radio propagation environment is significantly affected by various factors such as wind speed, temperature, sea states, and vessel motion. Additionally, scatterers and users are usually sparsely distributed over the vast sea surface, rendering traditional assumptions like Rayleigh fading invalid. Instead, maritime wireless channels exhibit sparsity and location dependency[3]. Furthermore, sea surface fluctuations are time-varying and not confined to fixed temporal scales, leading to more complex random undulations in nearshore wireless channels [15]. Thus, accurate modeling must account for the stochastic effects of wave dynamics.
@@
Among the existing maritime channel measurements, researchers mainly focused on the S-band and C-band. Lee et al. conducted land-to-ship channel measurements in the Yellow Sea of China, analyzing the received signal strength and the small-scale fading in the line-of-sight (LoS) propagation region at 2.4 GHz, and the results indicated that small-scale fading is best characterized by the Rician distribution [16]. Authors in [17] investigated the applicability of several terrestrial wireless channel models in maritime environments at 2 GHz and found that the ITU-R P.1546-2 model[18] showed the best agreement with the measured results. By adding a Gaussian random variable to the length of the rays in the traditional two-ray model, the work [19] described the randomness introduced by environmental changes in maritime channels. The researchers in [20] conducted land-to-ship channel measurements at 3.2 GHz and analyzed how islands affect wireless propagation, revealing that cluster delays shifted linearly with distance and small-scale fading changed from Rician to log-normal distribution when the ship passed through islands. In [21], measurements were conducted at 5 GHz in tropical sea regions, and it was proposed that a third refracted path, in addition to the direct and reflected paths, may occur during long-distance (beyond the break-point distance in the two-ray model) maritime transmission due to the evaporation duct effect over the sea. The measurement results showed excellent agreement with the theoretical curves [21], indicating that considering the duct effect is necessary in long-distance maritime scenarios. In[22], channel measurements were carried out at 5.8 GHz and instantaneous power delay profiles (PDPs) of wireless links were acquired between buoys and ships under various multipath conditions, followed by the investigation on delay dispersion characteristics. In [23], 5.8 GHz fixed WiMAX measurements were conducted in harbor environments to study the impacts of multipath effects, Doppler shift, and vessel motion on wireless channels. For the same frequency band, a land-to-ship channel measurement campaign was launched [24, 25], accompanied with some preliminary analysis on path loss, small-scale fading, and root-mean-square (RMS) delay spread. Ship-to-ship measurements at 5.9 GHz with a maximum distance of 2.6 km were performed[26], where the statistical properties of several channel parameters were analyzed. The authors of [27] conducted measurements with different receiver (Rx) heights at 8 GHz in the South China Sea. Among the fitting results of various path loss models, the log-distance model exhibited the smallest root mean square error (RMSE), while the parabolic equation model provided additional information on interference nulls. For small-scale fading, the Rician distribution showed the best Kolmogorov–Smirnov (K-S) test metrics compared to other commonly used distributions.
@@
Attention is also gradually shifting towards even higher frequency bands. A beyond-the-horizon maritime channel measurement campaign was conducted at 25 GHz for fixed-to-fixed point links[28]. Based on the measurement data, the authors analyzed the short-term variations in received signal power caused by environmental changes and presented empirical models. To better align with real-world conditions, some studies have applied modifications to the traditional two-ray model to more comprehensively account for fading produced by maritime environmental factors. For instance, [29] combined reflection from the sea surface and antenna height variation to propose a modified two-ray model through theoretical analysis, while [30] introduced an exponential coefficient into the two-ray model formula to correct for fluctuations affected by environmental factors via ray tracing, and achieved better prediction results in the millimeter-wave frequency band.
@@
To date, the existing works on maritime channel measurements and modeling have not considered the impact of various environmental factors on channel characteristics, or the investigated distances are limited, or the channel parameters and models are not sufficiently comprehensive. In this paper, real-world channel measurement data at 5.8 GHz for the land-to-ship scenario in the Yellow Sea of China [24, 25] is used to conduct systematic modeling and analysis on a variety of key channel parameters such as large-scale path loss, power fading on different time scales, channel sparsity, and temporal statistics. The main novelty and contributions of this work, compared with the current literature including our previous studies [24] and [25], are as follows:
@@
  • A series of large-scale fading models are employed to characterize the location-dependent path loss. More importantly, a new dual-slope large-scale path loss model is propounded, which embodies a solid physical basis, high modeling accuracy, and super adaptability over a wide range of distances and sea conditions. In addition, the two-wave with diffuse power (TWDP) model that has Rayleigh and Rician distributions as special cases [31] and asymmetric Laplace function are utilized to depict the small-scale fading, in addition to the conventional Rician, Nakagami-m, and lognormal distributions.
@@
  • Channel sparsity for the marine propagation environment is investigated in depth, where different metrics including the Gini index and Rician $K$ factor are adopted to quantify the sparsity level. The Gini index stems from economics [32] and has been shown to predict channel sparsity accurately, and it is garnering attention recently in research on land channels but has not been considered for maritime scenarios so far to our best knowledge.
The remainder of this paper is organized as follows. In Section II, the maritime wireless channel measurement system and procedure are presented, along with a brief description of the data processing method. Sections III to V focus on the large-scale, SWIFT, and small-scale fading, respectively. Channel sparsity is analyzed in Section VI, followed by the temporal statistics in Section VII. Conclusions are drawn in Section VIII.
@@ II-A Measurement System
At the Rx, also shown in Fig. 1, the signal is received by an omnidirectional antenna with a gain of 13 dBi. The received signal is then amplified by a low noise amplifier with 27.4 dB gain and subsequently processed by a spectrum and signal analyzer (R&S® FSW43) for analog-to-digital conversion and IQ demodulation. A 10 MHz GPS signal is employed at both the Tx and Rx to synchronize their respective transmission and reception times. Environmental information is collected along with the measured data. Fig. 2 presents photographs of the measurement equipment and the overall measurement scenario over the open sea. More details of the measurement system parameter settings are listed in Table I.
@@ II-B Measurement Procedure
The Tx antenna was fixed on the rooftop of a 25 m-high residential building, while the Rx antenna was mounted on the rail of a vessel’s deck as shown in Fig. 2(c). A Kestrel 5500 weather meter was installed on the vessel to record meteorological parameters such as the wind speed, air temperature, humidity, and atmospheric pressure, as displayed in Fig. 2(d). Additional marine environmental information, including sea water temperature and wave height, was recorded using the monitoring buoys in the sea area of Qidong City, as illustrated in Fig. 2(e). The measurements were carried out over two consecutive days. On the first day, the average relative wind speed over the sea surface was about 7.7 m/s, higher than the average relative wind speed of 5.6 m/s on the second day, while the other measurement conditions were almost the same. Correspondingly, there were significant differences in the sea wave conditions between the two days.
[…]
@@ II-C Data Processing
where $\otimes$ and * denote the correlation and convolution operations, respectively, $h(t,\tau)$ is the ground-truth CIR. The PDP is the squared-magnitude of the CIR, i.e., $|\hat{h}(t,\tau)|^{2}$. Examples of measured PDPs in the continuous-distance measurements are portrayed in Fig. 4.  [^3] It is noteworthy that the received power drops substantially around 4 km compared with its adjacent distances in Fig. 4, corresponding to the most dominant peak in path loss in Fig. 5(b), which is discussed in Section III.
@@
A comprehensive characterization of maritime channels requires an understanding of different fading mechanisms occurring across distinct temporal scales. The instantaneous received power can be conceptually expressed as
@@ III-A Empirical Distance-Dependent Path Loss Models
$d_{\text{LoS}}=\sqrt{h_{\text{t}}^{2}+2h_{\text{t}}r_{\text{e}}}+\sqrt{h_{\text{r}}^{2}+2h_{\text{r}}r_{\text{e}}},$ (4)
@@
$\text{PL}^{\text{two-ray}}[\text{dB}]=20\log_{10}\left[\frac{4\pi d}{\lambda}\bigg/\left|2\sin\left(\frac{2\pi h_{\text{t}}h_{\text{r}}}{\lambda d}\right)\right|\right],$ (5)
@@
In realistic maritime channels, unlike the traditional two-ray model, which assumes an idealized reflection coefficient of -1, we develop a modified two-ray (MTR) model influenced by multiple physical factors as follows
$\text{PL}^{\text{MTR}}[\text{dB}]=20\log_{10}\left[\frac{4\pi d}{\lambda\left|1+DSR\Gamma\exp\left(-j\frac{2\pi\Delta d}{\lambda}\right)\right|}\right],$ (6)
@@
$D=\left[1+\frac{2d_{1}d_{2}}{r_{e}(h_{\text{t}}+h_{\text{r}})}\right]^{-\frac{1}{2}},$ (7)
@@
$\Lambda(\tan\theta_{i})=\frac{1}{2}\Bigg\{\sqrt{\frac{2}{\pi}}\frac{\beta_{0}\exp[{-\left(\frac{\tan\theta_{i}}{\sqrt{2}\beta_{0}}\right)^{2}}]}{\tan\theta_{i}}-\text{erfc}\left(\frac{\tan\theta_{i}}{\sqrt{2}\beta_{0}}\right)\Bigg\}.$ (9)
@@
Figure 5: Measured data and various path loss models for the two measurement days, with parameters and performance metrics summarized in Table III.
@@
In (12) and (13), $\text{PL}^{\text{CI}}$ and $\text{PL}^{\text{Dual-CI}}$ denote the path loss in dB over frequency and distance, $n$, $n_{1}$, and $n_{2}$ represent the path loss exponents (PLEs) within different distance ranges, and $X_{\sigma}^{\text{CI}}$ and $X_{\sigma}^{\text{Dual-CI}}$ are zero-mean log-normally distributed random variables describing large-scale shadow fading (SF).
@@ III-B Measured Results and Proposed Path Loss Model
$\color[rgb]{0,0,0}\text{PL}^{\text{Dual-CI-MTR}}(f,d)[\text{dB}]\\ =\begin{cases}10n_{1}\log_{10}\left[\frac{4\pi fd}{c\left|1+DSR\Gamma\exp\left(-j\frac{2\pi f\Delta d}{c}\right)\right|}\right],\quad\quad\quad\quad\quad\quad\quad\quad~\text{for }d\leq d_{\text{break}}\\ 10n_{1}\log_{10}\left[\frac{4\pi fd_{\text{break}}}{c\left|1+DSR\Gamma\exp\left(-j\frac{2\pi f\Delta d}{c}\right)\right|}\right]+10n_{2}\log_{10}\left(\frac{d}{d_{\text{break}}}\right),\text{ }\text{for }d>d_{\text{break}}\end{cases}+X_{\sigma}^{\text{Dual-CI-MTR}}$} (14)
@@
In view of the trend of the measured data and observations in Fig. 5, we propound a new path loss model named dual-slope CI-MTR model integrating the CI and MTR models, which is expressed in (14). In this model, the MTR model incorporating an adaptive PLE $n_{1}$ is proposed to delineate the path loss within $d_{\text{break}}$, in conjunction with a second slope $n_{2}$ similar to that in the dual-slope CI model for distances beyond $d_{\text{break}}$. As listed in Table III, the slope values in the dual-slope CI-MTR model are similar to those in the dual-slope CI model in relevant literature (e.g., around 2 and 6 for calm sea in[23]), with lower $\sigma$ (e.g., over 8 dB in the literature) probably due to improved fitting within $d_{\text{break}}$. Since no standardized maritime path loss model with PLEs exists to our best knowledge, we compare the PLEs and SF standard deviations in our dual-slope CI-MTR model against 3GPP TR 38.901’s rural macro (RMa) LoS scenario. Most PLEs align (2 and 4 for the two slopes), except for $n_{2}$ on Day 2, likely due to Earth-curvature diffraction effects absent in the 3GPP RMa’s 10 km range limit. Besides, the $\sigma$ on Day 1 is higher than the RMa scenario (4-6 dB), implying windy environments render more widely spread path loss values.
Remark 1: The dual-slope CI-MTR model captures the oscillating behavior of the path loss within $d_{\text{break}}$, as demonstrated by the multiple peaks of the dashed blue curve in Fig. 5. Furthermore, it allows for more flexible and accurate characterization of the path loss based on environmental conditions via $n_{1}$ and $n_{2}$.